# MAPGD: Multi-Agent Prompt Gradient Descent for Collaborative Prompt Optimization

### v2

@@ Appendix A CAAW Parameter Sensitivity
Table 4. Test accuracy of MAPGD under different CAAW weighting parameter $\lambda$. Performance differences are minor, demonstrating robustness to $\lambda$.
@@
Appendix B Complexity and Parallelism
@@
Algorithm 1 MAPGD with HCGC and CAAW
3:for $t=1$ to $R$ do
4:  $M_{t}\leftarrow\textsc{SampleMiniBatch}(D_{\text{train}},b)$
5:  $\mathcal{G}_{t}\leftarrow\textsc{GenerateAgentGradients}(\{A_{i}\},M_{t})$ $\triangleright$ Alg. 2
7:  $F_{t}\leftarrow\textsc{CAAWFuse}(\tilde{\mathcal{G}}_{t},D_{\text{dev}})$ $\triangleright$ Alg. 4
[…]
12:  end if
13:end for
14:return Best prompt over $\{p_{\star}^{(1)},\dots,p_{\star}^{(t)}\}$
@@
Appendix C Theoretical Analysis
@@ C.3. Main Results
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\!\left[\|\nabla F(p^{(t)})\|^{2}\right]\;\leq\;\frac{2(F(p^{(1)})-F_{\inf})}{\mu T\eta}+\frac{L\sigma^{2}}{\mu}\eta.$
@@
C.4. Summary of Theoretical Guarantees
@@
  • Directional Alignment (A1): Although textual pseudo-gradients are not true mathematical gradients, Hypersphere-Constrained Gradient Clustering aggregates semantically coherent suggestions from multiple agents into a fused gradient that statistically aligns with the true descent direction.
@@
Appendix D Case Study: System Prompt Optimization
@@
Appendix E Agent Specialization for Different Tasks
@@
E.1. Agent Roles for Classification Tasks
@@
E.2. Agent Roles for Mathematical Reasoning Tasks
@@
E.3. Agent Roles for Financial Report Analysis
@@
Appendix F Example Prompts Before and After Optimization
@@
F.1. LIAR Dataset
Initial Prompt
# Task
Determine whether the Statement is a lie (Yes) or not (No)
based on the Context and other information.
# Output format
Answer Yes or No as labels
# Prediction
Text: {text}
label:
Optimized Prompt
# Task
Your task is to evaluate the given Statement in the context
of the provided information. If the Statement is found to be
intentionally false or misleading, it should be classified
as a lie. In this case, your response should be ’Yes’.
If the Statement is true or unintentionally false due to a
mistake or lack of knowledge, it should not be classified
as a lie. In this case, your response should be ’No’.
# Output format
Provide your answer as ’Yes’ or ’No’ labels. ’Yes’ indicates
that the statement is a lie, while ’No’ indicates that the
statement is not a lie.
# Prediction
Text: {text}
label:
F.2. Ethos Dataset
Initial Prompt
# Task
Is the following text hate speech?
# Output format
Answer Yes or No as labels
# Prediction
Text: {text}
Label:
Optimized Prompt
# Task
Your task is to analyze the following text and determine if
it contains hate speech. For the purpose of this task, hate
speech is defined as any speech, gesture, conduct, writing,
or display that may incite violence or prejudicial actions
against or by a protected individual or group, or because it
disparages or intimidates a protected individual or group.
Please base your judgment on the explicit content of the text,
rather than personal interpretations or assumptions.
# Output format
Provide your answer in a clear format as either
’Contains Hate Speech’ or ’Does Not Contain Hate Speech’
# Prediction
Text: {text}
Label:
F.3. Jailbreak Dataset
Initial Prompt
# Task
Detect if the message is a jailbreak attack,
i.e. an attempt by a user to break through an
AI system’s protections
# Output format
Answer Yes or No as labels
# Prediction
Text: {text}
Label:
Optimized Prompt
# Task
Classify the message as one of the following labels based on
its potential to be a jailbreak attack, i.e., an attempt to
bypass an AI system’s protections:
- "OvertAttack": The message contains clear, direct attempts
  to break through system safeguards.
- "SuspiciousContext": The message shows indirect or
  contextual cues that may indicate a multi-stage or
  sophisticated attack attempt.
- "Benign": The message poses no apparent risk and does not
  attempt to circumvent protections.
# Output format
Select and output exactly one label from the options above.
# Prediction
Text: {text}
Label:
F.4. GSM8K Dataset
Initial Prompt
# Task
Solve the math word problem step by step.
# Instructions
1. Read the problem carefully
2. Identify what needs to be calculated
3. Show your work step by step
4. Provide the final numerical answer
# Output Format
Show your reasoning process and end with: #### [final_answer]
# Problem
{text}
Optimized Prompt
# Task
Solve the math word problem step by step.
# Instructions
1. Read the problem carefully
2. Identify what needs to be calculated
3. If there are ambiguous situations in the problem, make
   reasonable assumptions and state them clearly
4. Show your work step by step
5. Provide the final numerical answer
# Output Format
Show your reasoning process, including any assumptions made,
and end with: #### [final_answer]
# Problem
{text}
F.5. AQUARAT Dataset
Initial Prompt
# Task
Solve the math word problem and choose the correct answer
from the given options.
# Instructions
1. Read the problem carefully
2. Analyze each option
3. Show your reasoning step by step
4. Select the correct answer (A, B, C, D, or E)
# Output Format
Show your reasoning and end with: Answer: [LETTER]
(For example: "Answer: A" or "Answer: B")
# Problem
{text}
# Options
{options}
Optimized Prompt
# Task
Solve the provided problem by systematically classifying its
type, analyzing all constraints, applying formal mathematical
reasoning, and verifying the solution against all conditions
before selecting the correct answer.
# Instructions
Adhere strictly to the following structured framework:
1. **Problem Classification:**
   * Explicitly identify the core problem domain (e.g.,
     "Combinatorial Selection," "Constraint Satisfaction,"
     "Partnership Profit-Sharing," "Work Rate").
   * Justify the classification with a brief rationale
     based on the problem statement.
2. **Constraint Analysis:**
   * Enumerate all explicit constraints from the problem text.
   * Infer and list any implicit constraints or logical
     dependencies.
   * Categorize each constraint (e.g., Hard/Must-Fulfill,
     Soft/Optimization) and specify their logical
     relationships (e.g., AND, OR).
3. **Mathematical Formalization & Numerical Context
     Resolution:**
   * Identify and resolve any ambiguous numerical references
     (e.g., "monthly" vs. "annual", "together" vs.
     "individual").
   * Infer and incorporate any implicit numerical variables
     (e.g., initial quantities, unstated rates, starting
     points).
   * Normalize all units to a consistent basis for calculation.
   * Define all variables, sets, and parameters using clear
     notation (e.g., C_total, P_A, S_eligible, nCr).
   * Structure the problem using appropriate formalisms:
     timelines, sets, equations, or logical statements.
   * For combinatorial problems, explicitly state
     the formula used.
4. **Solution Execution:**
   * Perform all calculations step-by-step, showing
     substitutions and operations.
   * For combinatorial scenarios, enumerate valid
     combinations or calculate cardinalities.
   * Derive the target value (e.g., profit share, number of
     committees, optimal value).
5. **Verification & Solution Finalization:**
   * Systematically verify that the proposed solution
     satisfies every constraint from Step 2.
   * For combinatorial answers, confirm counts against
     constraints using direct checks or complementary
     counting.
   * Match the final, verified result to the provided
     options to select the correct answer.
# Output Format
Your final output must follow this structure precisely:
**Problem Classification:**
[Your classification and rationale]
**Constraint
F.6. SVAMP Dataset
Initial Prompt
# Task
Solve the math word problem step by step.
# Instructions
1. Read the problem carefully
2. Identify what needs to be calculated
3. Show your work step by step
4. Provide the final numerical answer
# Output Format
Show your reasoning process and end with: #### [final_answer]
# Problem
{text}
Optimized Prompt
F.7. SVAMP Dataset
Optimized Prompt
# Task
Solve the math word problem through structured, hierarchical
reasoning. Your primary goal is to correctly parse and compute
multi-step problems involving distinct object categories and
their quantitative relationships.
# Instructions
1. **Parse and Categorize:**
   Read the problem and question carefully. Identify and list
   all distinct object types or categories (e.g., ’packages
   of gum’, ’boxes of candy’). For each category, extract:
   - The number of units (e.g., 5 packages).
   - The quantity per unit, if specified (e.g., 8 pieces
     per package).
2. **Structure the Solution Hierarchically:**
   - **Step 1: Calculate Intra-Category Totals.**
     For each identified category, calculate its total quantity.
     If a unit rate is given, perform the multiplication
     (e.g., ‘5 packages * 8 pieces/package = 40 pieces of gum‘).
     If no unit rate exists, use the given quantity directly.
   - **Step 2: Perform Inter-Category Operations.**
     Using the totals from Step 1, now perform the final operation
     as required by the question (e.g., sum all category totals
     for a grand total, or find the difference between two
     category totals for a comparison).
3. **Show Your Work:**
   Present your reasoning clearly, reflecting this two-step
   hierarchical process. First, show all calculations within
   categories. Then, show the final combination of these
   category totals.
4. **Final Answer:**
   Box your final numerical result.
# Output Format
Reasoning: [Your step-by-step reasoning here]
#### [final_answer]
# Problem
{text}
# Question
{question}

### v3

@@ Appendix D CAAW Parameter Sensitivity
Table 7: Test accuracy of MAPGD under different CAAW weighting parameter $\lambda$. Performance differences are minor, demonstrating robustness to $\lambda$.
@@
Appendix E Comparison with Recent Concurrent Work
We are aware of two recent concurrent works that are closely related in spirit to prompt optimization with multi-agent components, namely MAPRO Zhang et al. (2025) and MA-SAPO Seo et al. (2025). While these works share a high-level motivation with MAPGD, we did not include them as baselines in our main experimental comparisons due to comparability constraints in problem setting and resource assumptions.
MAPRO Zhang et al. (2025).
MAPRO studies prompt optimization within an explicit multi-agent system (MAS): it jointly configures role prompts for multiple agents and leverages the MAS interaction topology (e.g., DAG-structured communication) together with node-/edge-level reward modeling and global inference to select prompt combinations. In contrast, MAPGD targets task-level prompt optimization under a fixed query budget, where specialized agents propose pseudo-gradients that are subsequently coordinated (via HCGC and CAAW) to iteratively refine a single task prompt. Because MAPRO’s formulation depends on an explicit MAS graph, per-agent prompts, and interaction-specific rewards, directly porting it to our benchmark protocol would require redesigning the environment and supervision signals, making a head-to-head comparison not directly aligned with our evaluation setting.
MA-SAPO Seo et al. (2025).
MA-SAPO optimizes prompts through retrieved reasoning assets: it constructs a searchable library of semi-structured “reasoning cards” from labeled or scored prompt–response data, and then performs analysis–refinement guided by top-$k$ retrieved assets at test time. This introduces an additional external memory/resource assumption (i.e., an offline asset corpus and its construction pipeline) that is orthogonal to our training-free, black-box optimization protocol. In our evaluation, we intentionally restrict methods to operate under the same query budget without relying on extra annotated corpora or a separately curated retrieval memory. Under these constraints, implementing MA-SAPO would either (i) require additional data collection/annotation to build assets, or (ii) reuse task-specific optimization traces as assets, which changes the method definition and complicates fairness.
Summary and future work.
Overall, MAPRO and MA-SAPO represent complementary directions—MAS-level joint role-prompt configuration and retrieval-augmented prompt rewriting—whereas MAPGD focuses on multi-agent pseudo-gradient coordination for budgeted task-level prompt refinement. Integrating these paradigms into a unified and strictly comparable evaluation (e.g., extending MAPGD with optional retrieval memories or evaluating MAPRO-style joint prompting on explicit MAS benchmarks) is an important direction for future work.
Appendix F Complexity and Parallelism
@@
Appendix G Theoretical Analysis
@@ G.3 Main Results
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\!\left[\|\nabla F(p^{(t)})\|^{2}\right]\;\leq$ $\frac{2\big(F(p^{(1)})-F_{\inf}\big)}{\mu T\eta}$
$\;+\;\frac{L\sigma^{2}}{\mu}\eta.$
@@
G.4 Summary of Theoretical Guarantees
@@
  • Directional Alignment (A1): Although textual pseudo-gradients are not true mathematical gradients, HCGC aggregates semantically coherent suggestions from multiple agents into a fused gradient that statistically aligns with the true descent direction.
@@
Appendix H Qualitative Case Study: From Conflicting Pseudo-Gradients to Prompt Updates
H.1 Case Setup
We select a representative optimization step on the LIAR task where multiple specialized agents propose semantically conflicting pseudo-gradients. This step exhibits clear disagreement among agents while still leading to a measurable improvement in downstream validation behavior after gradient fusion.
Item Value
Task LIAR
Iteration $t=3$
Agents 4
Raw Gradients 16
Illustrated Conflict Pairs 2
HCGC Clusters 3
At this iteration, agents disagree on whether errors primarily stem from ambiguous label semantics, missing contextual cues, or insufficient decision boundary clarification, making it a suitable case for qualitative inspection.
H.2 Raw Pseudo-Gradients from Specialized Agents
We focus on two representative conflict pairs extracted from the same optimization step, which illustrate structurally different failure diagnoses and motivate the need for explicit conflict handling.
Conflict Pair 1: Format Clarity vs Example Diversity.
Format Designer:
<START> The prompt instructs the model to answer with ‘Yes’ or ‘No’, but it does not specify what these labels represent. This ambiguity can cause the model to misinterpret the classification objective. <END>
Example Curator:
<START> The provided examples do not cover a sufficiently diverse range of statements. Including more varied and representative cases could help the model generalize better. <END>
The cosine similarity between these two gradients is $0.13$ (angle =1.44=1.44 rad), indicating strong semantic disagreement. Importantly, the conflict is not superficial: the format-oriented gradient advocates tightening the input–output contract through explicit label semantics, while the example-oriented gradient pushes the prompt toward structural expansion via additional demonstrations. Applying both updates independently would lead to incompatible prompt modifications—one emphasizing deterministic label interpretation and the other reallocating prompt capacity to example coverage.
Conflict Pair 2: Example Coverage vs Definition of “Lie”.
Example Curator:
<START> The examples fail to illustrate borderline or ambiguous cases, which may limit the model’s ability to generalize across different claim types. <END>
Instruction Specialist:
<START> The prompt does not clearly define what constitutes a ‘lie’ versus an inaccurate or exaggerated statement, leading to inconsistent decisions. <END>
This pair exhibits a cosine similarity of $0.23$ (angle =1.34=1.34 rad), reflecting a fundamental disagreement about the source of errors. The example-focused gradient assumes that misclassification primarily stems from insufficient empirical coverage, whereas the instruction- focused gradient attributes failures to an underspecified decision criterion. These two perspectives induce different optimization directions: expanding illustrative evidence versus formalizing the semantic boundary of the task itself.
Discussion.
Both conflict pairs arise from valid but incompatible diagnostic assumptions. Crucially, neither gradient is inherently incorrect; each targets a distinct aspect of model behavior. This highlights the core challenge addressed by MAPGD: without explicit conflict detection and coordination, naïvely aggregating such gradients would obscure their semantic tension and lead to unstable or diluted prompt updates.
H.3 Conflict Detection and HCGC Clustering
Based on the two conflict pairs in §H.2, we observe clear semantic disagreement among agent-proposed directions. In particular, the Format Clarity vs Example Diversity pair has cosine similarity $0.13$ (angle =1.44=1.44 rad), and the Example Coverage vs Definition of “Lie” pair has cosine similarity $0.23$ (angle =1.34=1.34 rad). The corresponding angles ($1.44$ and $1.34$ rad) indicate substantial directional mismatch, and both similarities fall below the conflict threshold. HCGC therefore marks them as conflicting and avoids naïve averaging.
After projecting all gradient embeddings onto the unit hypersphere and applying the angular-margin constraints, HCGC partitions the 16 raw pseudo-gradients into three coherent clusters, as summarized in Table 8.
Table 8: HCGC clustering summary for the representative LIAR optimization step (t=3t{=}3).
Cluster Agents Involved Semantic Theme
C1 instruction, format label semantics and task definition (binding Yes/No to True/False, clarifying what is being judged)
C2 example, style coverage and contextual cues for borderline cases (exaggeration, unverifiable framing, pragmatic signals)
C3 instruction only decision-boundary calibration (strictness vs recall; discouraging overly permissive acceptance of absolute claims)
This clustering reflects a key property of HCGC: conflicting gradients are not forced into a single averaged direction. Instead, they are separated into semantically consistent groups (Table 8), enabling subsequent fusion to preserve complementary information (e.g., clarifying label semantics in C1 while retaining coverage-oriented refinements in C2) without letting one channel dominate the other. The semantic themes are manually summarized by the authors to reflect the dominant intent of each cluster, rather than being directly emitted by the model.
H.4 Channel-Adaptive Agent Weighting
Given the clustered directions above, CAAW assigns each agent a contribution weight based on its recent performance signal, so that the subsequent fusion emphasizes gradients that are empirically reliable for the current task state. At this iteration, the adaptive weights are:
[…]
These weights are consistent with the observed conflict structure. Clusters C1 and C2 correspond to two structurally competing but both valid refinement axes: (i) clarifying label semantics and the task definition (C1), and (ii) improving coverage and contextual handling of borderline cases (C2). CAAW prioritizes instruction- and example-driven gradients while down-weighting stylistic adjustments, preventing prompt capacity from being consumed by surface-level phrasing when the dominant failures are semantic (label meaning) and epistemic (what constitutes a lie).
H.5 Fused Gradient and Prompt Modification
After separating conflicting directions with HCGC and weighting them with CAAW, MAPGD fuses the cluster-level signals into a single coherent update that preserves complementary intent across clusters. In this case, the fused direction primarily integrates (a) C1: explicit label semantics and task definition, and (b) C2: guidance for borderline or unverifiable claims, while avoiding excessive prompt expansion.
Before:
“The following statement is either true or false. Answer Yes or No.”
After:
“Determine whether the statement is factually accurate. Treat Yes as True and No as False. If the statement relies on exaggeration, satire, or unverifiable claims, judge it as False. Return only Yes or No.”
This modification reflects the fused gradient in two concrete ways: (i) it resolves the format/label ambiguity highlighted in C1 by explicitly binding Yes/No to True/False, and (ii) it incorporates the borderline-case handling emphasized by C2 without requiring lengthy demonstrations, thereby avoiding the prompt-structure conflict that would arise from naïvely mixing rule-heavy and example-heavy updates.
H.6 Behavioral Change on Previously Misclassified Examples
We report qualitative behavioral changes on examples that were previously misclassified due to (i) label ambiguity and (ii) overly permissive interpretation of absolute or exaggerated claims.
Example 1 (Absolute claim / overgeneralization).
Claim: “The senator has voted against every environmental bill.”
Before: Yes   After: No
The updated prompt instructs the model to treat unverifiable or sweeping claims as False, which directly corrects the previous tendency to accept absolute statements (e.g., “every”) as True.
Example 2 (Borderline / unverifiable framing).
Claim: “The policy has been a complete disaster for all families.”
Before: Yes   After: No
This change follows the newly added rule that exaggeration or unverifiable claims should be judged as False, reducing false positives driven by rhetorical or hyperbolic language (e.g., “complete disaster”).
Appendix I Case Study: System Prompt Optimization
@@
Appendix J Agent Specialization for Different Tasks
@@
J.1 Agent Roles for Classification Tasks
@@
J.2 Agent Roles for Mathematical Reasoning Tasks
@@
J.3 Agent Roles for Financial Report Analysis
@@
Appendix K Example Prompts Before and After Optimization
@@
K.1 LIAR Dataset
Initial Prompt # Task Determine whether the Statement is a lie (Yes) or not (No) based on the Context and other information. # Output format Answer Yes or No as labels # Prediction Text: {text} label:
Optimized Prompt # Task Evaluate the given Statement using the provided context. Classify as Yes if the Statement is intentionally false or misleading (i.e., a lie). Classify as No if the Statement is true, or if it is incorrect due to a mistake or lack of knowledge rather than intentional deception. # Output format Output exactly one label: Yes or No’. Yes indicates the Statement is a lie; No indicates it is not a lie. # Prediction Text: {text} label:
K.2 Ethos Dataset
Initial Prompt # Task Is the following text hate speech? # Output format Answer Yes or No as labels # Prediction Text: {text} Label:
Optimized Prompt # Task Your task is to analyze the following text and determine if it contains hate speech. For the purpose of this task, hate speech is defined as any speech, gesture, conduct, writing, or display that may incite violence or prejudicial actions against or by a protected individual or group, or because it disparages or intimidates a protected individual or group. Please base your judgment on the explicit content of the text, rather than personal interpretations or assumptions. # Output format Provide your answer in a clear format as either Contains Hate Speech or Does Not Contain Hate Speech # Prediction Text: {text} Label:
K.3 Jailbreak Dataset
Initial Prompt # Task Detect if the message is a jailbreak attack, i.e. an attempt by a user to break through an AI systems protections # Output format Answer Yes or No as labels # Prediction Text: {text} Label:
Optimized Prompt # Task Classify the message as one of the following labels based on its potential to be a jailbreak attack, i.e., an attempt to bypass an AI systems protections: - "OvertAttack": The message contains clear, direct attempts to break through system safeguards. - "SuspiciousContext": The message shows indirect or contextual cues that may indicate a multi-stage or sophisticated attack attempt. - "Benign": The message poses no apparent risk and does not attempt to circumvent protections. # Output format Select and output exactly one label from the options above. # Prediction Text: {text} Label:
K.4 GSM8K Dataset
Initial Prompt # Task Solve the math word problem step by step. # Instructions 1. Read the problem carefully 2. Identify what needs to be calculated 3. Show your work step by step 4. Provide the final numerical answer # Output Format Show your reasoning process and end with: #### [final_answer] # Problem {text}
Optimized Prompt # Task Solve the math word problem step by step. # Instructions 1. Read the problem carefully 2. Identify what needs to be calculated 3. If there are ambiguous situations in the problem, make reasonable assumptions and state them clearly 4. Show your work step by step 5. Provide the final numerical answer # Output Format Show your reasoning process, including any assumptions made, and end with: #### [final_answer] # Problem {text}
K.5 AQUARAT Dataset
Initial Prompt # Task Solve the math word problem and choose the correct answer from the given options. # Instructions 1. Read the problem carefully 2. Analyze each option 3. Show your reasoning step by step 4. Select the correct answer (A, B, C, D, or E) # Output Format Show your reasoning and end with: Answer: [LETTER] (For example: "Answer: A" or "Answer: B") # Problem {text} # Options {options}
Optimized Prompt # Task Solve the provided problem by systematically classifying its type, analyzing all constraints, applying formal mathematical reasoning, and verifying the solution against all conditions before selecting the correct answer. # Instructions Adhere strictly to the following structured framework: 1. **Problem Classification:** * Explicitly identify the core problem domain (e.g., "Combinatorial Selection," "Constraint Satisfaction," "Partnership Profit-Sharing," "Work Rate"). * Justify the classification with a brief rationale based on the problem statement. 2. **Constraint Analysis:** * Enumerate all explicit constraints from the problem text. * Infer and list any implicit constraints or logical dependencies. * Categorize each constraint (e.g., Hard/Must-Fulfill, Soft/Optimization) and specify their logical relationships (e.g., AND, OR). 3. **Mathematical Formalization & Numerical Context Resolution:** * Identify and resolve any ambiguous numerical references (e.g., "monthly" vs. "annual", "together" vs. "individual"). * Infer and incorporate any implicit numerical variables (e.g., initial quantities, unstated rates, starting points). * Normalize all units to a consistent basis for calculation. * Define all variables, sets, and parameters using clear notation (e.g., C_total, P_A, S_eligible, nCr). * Structure the problem using appropriate formalisms: timelines, sets, equations, or logical statements. * For combinatorial problems, explicitly state the formula used. 4. **Solution Execution:** * Perform all calculations step-by-step, showing substitutions and operations. * For combinatorial scenarios, enumerate valid combinations or calculate cardinalities. * Derive the target value (e.g., profit share, number of committees, optimal value). 5. **Verification & Solution Finalization:** * Systematically verify that the proposed solution satisfies every constraint from Step 2. * For combinatorial answers, confirm counts against constraints using direct checks or complementary counting. * Match the final, verified result to the provided options to select the correct answer. # Output Format Your final output must follow this structure precisely: **Problem Classification:** [Your classification and rationale]
K.6 SVAMP Dataset
Initial Prompt # Task Solve the math word problem step by step. # Instructions 1. Read the problem carefully 2. Identify what needs to be calculated 3. Show your work step by step 4. Provide the final numerical answer # Output Format Show your reasoning process and end with: #### [final_answer] # Problem {text}
Optimized Prompt # Task Solve the math word problem through structured, hierarchical reasoning. Your primary goal is to correctly parse and compute multi-step problems involving distinct object categories and their quantitative relationships. # Instructions 1. **Parse and Categorize:** Read the problem and question carefully. Identify and list all distinct object types or categories (e.g., packages of gum’, boxes of candy’). For each category, extract: - The number of units (e.g., 5 packages). - The quantity per unit, if specified (e.g., 8 pieces per package). 2. **Structure the Solution Hierarchically:** - **Step 1: Calculate Intra-Category Totals.** For each identified category, calculate its total quantity. If a unit rate is given, perform the multiplication (e.g., ‘5 packages * 8 pieces/package = 40 pieces of gum‘). If no unit rate exists, use the given quantity directly. - **Step 2: Perform Inter-Category Operations.** Using the totals from Step 1, now perform the final operation as required by the question (e.g., sum all category totals for a grand total, or find the difference between two category totals for a comparison). 3. **Show Your Work:** Present your reasoning clearly, reflecting this two-step hierarchical process. First, show all calculations within categories. Then, show the final combination of these category totals. 4. **Final Answer:** Box your final numerical result. # Output Format Reasoning: [Your step-by-step reasoning here] #### [final_answer] # Problem {text} # Question {question}